OpenStudy (anonymous):

Simplify. square root of 5 over cube root of 5

OpenStudy (anonymous):

@precal @jdoe0001

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \frac{ \sqrt{5} } {\sqrt[3]{5}}= \frac{ \sqrt[2]{5} } {\sqrt[3]{5}}=\frac{5^{1/2}}{5^{1/3}} =5^{1/2-1/3} =5^{3/6-2/6} =5^{1/6} }\)

OpenStudy (solomonzelman):

You can re-write this as \(\huge\color{black}{ \bf \sqrt[6]{5} }\)

OpenStudy (anonymous):

Thanks! @SolomonZelman that makes so much sense!

OpenStudy (solomonzelman):

Anytime:) Good luck on your homework !

OpenStudy (anonymous):

Yea :)

OpenStudy (precal):

remember use your laws of exponents, this problem is just like the first one I helped you with